How many ways can three items be selected from a group of six items? Use the letters A,B,C,D,E, and F to identify the items, and list each of the different combinations of three items.
Answer by stanbon(75887)

How many ways can three items be selected from a group of six items? Use the letters A,B,C,D,E, and F to identify the items, and list each of the different combinations of three items.
Ans: 6C3 = (6*5*4)/(1*2*3) = 5*4 = 20
-------
I'll let you list them.
Cheers,
Stan H.
=================